# Geodesic completeness from string T-duality

### v1

@@ Abstract
1 Introduction
@@
To achieve our goals, we first introduce the string T-duality metric in Section 2. In Section 3, we study the Raychaudhuri equation for the string T-duality metric by presenting a careful analysis of the terms that determine the geodesic compression. In Section 4 we give our conclusions.
2 String T-duality metric
@@
[…]
@@
After these initial remarks, we introduce a modified Newtonian potential from the relation $f(r)=1+2\phi(r)$, with $\phi(r)=-m(r)/r$, where $m(r)$ is the cumulative mass distribution. From Fig. 1 we can see that $\phi(r)$ is attractive only at large distances, i.e. $r>1.7\ l_{0}$. Conversely, in a region of size about $l_{0}$ around the origin, quantum fluctuations dominate and their mean value locally forms a repulsive de Sitter core. The repulsive character is crucial for the following section, where the mechanism of singularity avoidance is analyzed in more detail.
@@
3 Geodesics in string T-duality metric
In this section we investigate the behavior of the expansion of geodesic congruence in the spacetime described by (2). For this purpose, the Raychaudhuri equation provides the rate of change of the expansion parameter, which in general depends on other quantities such as shear, twist and curvature.
@@ 3 Geodesics in string T-duality metric
Consider the normalized vector field $u^{a}$, namely the four-velocity in Painlevé–Gullstrand coordinates and a smooth congruence of time-like geodesics parametrized by the proper time $\tau$. Then the Raychaudhuri equation is
$\frac{d\Theta}{d\tau}=-\frac{1}{3}\Theta^{2}-\sigma_{\mu\nu}\sigma^{\mu\nu}+\omega_{\mu\nu}\omega^{\mu\nu}-R_{\mu\nu}u^{\mu}u^{\nu},\leavevmode\nobreak$ (6)
@@
$R_{\mu\nu}u^{\mu}u^{\nu}=8\pi\left(T_{\mu\nu}-\frac{1}{2}Tg_{\mu\nu}\right)u^{\mu}u^{\nu}=\rho+\sum_{i}P_{i}\geq 0,$ (15)
where $T_{\mu\nu}u^{\mu}u^{\nu}$ is the (positive definite) energy density of matter as measured by an observer at 4-velocity $u^{a}$, and $T=g_{\mu\nu}T^{\mu\nu}$ is the trace of the energy-momentum tensor ${T^{\mu}}_{\nu}=(-\rho,P_{r},P_{\theta},P_{\phi})$. [^1] The conservation of the energy-momentum tensor and the Schwarzschild-like condition $g_{00}=-1/g_{11}$ imply $P_{r}=-\rho$ and $P_{\theta}=P_{\phi}=-\rho-(r/2)\partial_{r}\rho$. The above equation is better known as the Strong Energy Condition (SEC). Its violation in the region $r<\sqrt{2/3}\,l_{0}$ is a piece of evidence for the non-classical nature of the spacetime geometry in the presence of T-duality effects.
@@
A closer look at (6) reveals other interesting aspects. First, (6) contains only even functions. This is due to the fact that the transformation $r\to-r$ produces a copy of the spacetime (2) rather than a negative-mass universe. Second, in the short-scale limit both $R_{\mu\nu}u^{\mu}u^{\nu}$ and $-\frac{1}{3}\Theta^{2}$ approach a constant value, while $\sigma_{\mu\nu}\sigma^{\mu\nu}$ vanishes rapidly, as already mentioned. However, the l.h.s. of (6) does not become constant in this limit. This can be explained by the fact that the contributions of $R_{\mu\nu}u^{\mu}u^{\nu}$ and $-\frac{1}{3}\Theta^{2}$ surprisingly cancel. Indeed, one finds
@@
Before proceeding, we recall that the present procedure can be extended to a generic cumulative mass distribution profile $m(r)$. Indeed, we find
@@
The case of time-like geodesics cannot change the conclusion that spacetime is geodesically complete. This can be inferred from the fact that the compression (9) and its evolution equation (16) depend on the properties of the background spacetime rather than on the nature of the propagating probe, namely massive or massless. Nevertheless, we briefly sketch the reasoning here.
@@
$\Box\longrightarrow\nabla_{\alpha}\nabla^{\alpha}-{R^{\nu}}_{\mu}$ (22)
where $\Box$ is the flat space d’Alembert operator and $R_{\mu\nu}\sim\Lambda_{\mathrm{eff}}\ \delta_{\mu\nu}$ is the Ricci tensor. This means that for photons the study of null geodesics becomes equivalent to that of time-like geodesics in the previous section.
@@ 4 Conclusions
Another important point is that the character of quantum gravity cannot depend on global properties of spacetime, such as symmetries and other features of the classical gravity regime. In fact, the attractive character of quantum gravity depends on the length scale $l_{0}$, i.e. on local spacetime properties. For this reason, our results also hold for other spacetime configurations, such as the state of the universe during the Planck epoch.
Finally, we emphasize that the above scenario is general and model independent. Alternative quantum gravity formulations may predict slightly different values of some parameters, but cannot change our conclusions: classical gravity gravitates, but quantum gravity antigravitates.
Acknowledgments

### v2

@@ Abstract
Singularity, stringy effects, quantum gravity
preprint: APS/123-QED
I Introduction
@@
II String T-duality metric
@@
Following the original derivation [27], one starts by determining the static potential due to the virtual particle exchange between the source $M$ and a generic probe mass. In doing so, one implements short scale quantum modifications using the T-duality modified propagator proposed by Padmanabahn, Spallucci and collaborators [35, 36, 37, 38]. The resulting Newton-Poisson equation no longer describes a point-like object. In fact, the mass density $\rho$ is distributed over a distance $l_{0}$, the so-called “zero point length”, with $l_{0}=2\pi\sqrt{\alpha^{\prime}}$ where $\alpha^{\prime}$ is the Regge slope. More importantly, the mass density is equal to $T_{0}^{\ 0}$, i.e. the $00$-component of the energy-momentum tensor, which goes into the r.h.s. of Einstein’s equations. So the solution is
@@
After these initial remarks, we introduce a modified Newtonian potential from the relation $f(r)=1+2\phi(r)$, with $\phi(r)=-m(r)/r$, where
$m(r)=\frac{Mr^{3}}{(r^{2}+l_{0}^{2})^{3/2}}$ (3)
is the cumulative mass distribution. From Fig. 1 we can see that $\phi(r)$ is attractive only at large distances, i.e. $r>1.7\ l_{0}$. Conversely, in a region of size about $l_{0}$ around the origin, quantum fluctuations dominate and their mean value locally forms a repulsive de Sitter core.
@@
III Geodesics in string T-duality metric
@@
Consider the normalized vector field $u^{a}=(1,-v(r),0,0)$, namely the four-velocity in Painlevé–Gullstrand coordinates and a smooth congruence of time-like geodesics parametrized by the proper time $\tau$. Then the Raychaudhuri equation is
$\frac{d\Theta}{d\tau}=-\frac{1}{3}\Theta^{2}-\sigma_{\mu\nu}\sigma^{\mu\nu}+\omega_{\mu\nu}\omega^{\mu\nu}-R_{\mu\nu}u^{\mu}u^{\nu},$~ (7)
@@
$R_{\mu\nu}u^{\mu}u^{\nu}=8\pi\left(T_{\mu\nu}-\frac{1}{2}Tg_{\mu\nu}\right)u^{\mu}u^{\nu}=\rho+\sum_{i}P_{i}\geq 0,$
where $T_{\mu\nu}u^{\mu}u^{\nu}$ is the (positive definite) energy density of matter as measured by an observer at 4-velocity $u^{a}$, and $T=g_{\mu\nu}T^{\mu\nu}$ is the trace of the energy-momentum tensor ${T^{\mu}}_{\nu}=(-\rho,P_{r},P_{\theta},P_{\phi})$. The above equation is better known as the Strong Energy Condition (SEC). Its violation in the region $r<\sqrt{\frac{2}{3}}l_{0}$ results from the presence of stringy effects encoded in the line elements (2). In principle, the violation of this condition could also occur in the absence of quantum effects, but for the rest of the paper, we will consider only the T-duality case or similar cases where the length scale at which there is a departure from general relativity arises from a quantum gravity formulation.
@@
A closer look at (7) reveals other interesting aspects. First, (7) contains only even functions. This is due to the fact that the transformation $r\to-r$ produces a copy of the spacetime (2) rather than a negative-mass universe. Second, in the short-scale limit both $R_{\mu\nu}u^{\mu}u^{\nu}$ and $\frac{1}{3}\Theta^{2}$ approach a constant value, while $\sigma_{\mu\nu}\sigma^{\mu\nu}$ vanishes rapidly, as already mentioned. However, the l.h.s. of (7) does not become constant in this limit. This can be explained by the fact that the contributions of $R_{\mu\nu}u^{\mu}u^{\nu}$ and $\frac{1}{3}\Theta^{2}$ surprisingly cancel. Indeed, one finds
@@
Before proceeding, we recall that the present procedure can be extended to a generic cumulative mass distribution profile $m(r)$ that satisfies regularity conditions at the origin, specifically $m(r)\sim r^{\alpha}$ with $\alpha\geq 3$ as $r\to 0$, and guarantees the existence of a proper definition of the ADM mass at infinity, namely $m(r)\to\text{constant}$ as $r\to\infty$. In such a case, we find
@@
It is essential to note that any non-Schwarzschild black hole metric exhibits a two-horizon configuration. In other words, the Schwarzschild geometry represents an unstable configuration in parameter space, where any modifications, such as charge or angular momentum, can trigger the transition to a two-horizon geometry. The “zero point length” parameter $l_{0}$ is no exception. However, the local and global behaviors of the manifold are independent concepts, and (22) captures only the local behavior. This can be easily verified by comparing the Reissner-Nordström geometry with the T-duality metric, both of which exhibit similar global behaviors (two horizons), yet their short-scale behaviors are drastically different, i.e., divergent versus regular.
The case of null geodesics cannot change the conclusion that spacetime is geodesically complete. This can be inferred from the fact that the compression (10) and its evolution equation (17) depend on the properties of the background spacetime rather than on the nature of the propagating probe, namely massive or massless. Nevertheless, we briefly sketch the reasoning here.
@@
$\Box\longrightarrow\nabla_{\alpha}\nabla^{\alpha}-R$ (23)
where $\Box$ is the flat space d’Alembert operator and $R=R_{\mu\nu}\ \delta_{\mu\nu}$, with $R_{\mu\nu}\sim\Lambda_{\mathrm{eff}}\ \delta_{\mu\nu}$ is the Ricci tensor. This means that for photons the study of null geodesics becomes equivalent to that of time-like geodesics.
@@ IV Conclusions
Another important point is that the character of quantum gravity cannot depend on the global properties of spacetime, such as symmetries and other features of the classical gravity regime. In fact, the attractive nature of quantum gravity depends on the length scale $l_{0}$, i.e., on local spacetime properties. This confirms that the genuine characteristic of quantum gravity is the intrinsic loss of resolution at the Planck scale, which remains unaffected regardless of the overall structure of the metric at large distances. An illustrative example can also be found in the regime of semiclassical gravity, where the conformal anomaly of quantum fields is a purely geometric local object arising from renormalization in curved space. It does not depend on quantum states over the manifold, which instead rely on global properties, such as the existence of a suitable coordinate covering [10, p. 178].
Turning to the specifics of our analysis, (22) does not rely on global properties, as it is sensitive to the short-distance behavior of the manifold. This sensitivity is highlighted by the scalar nature of the compression parameter $\Theta$. The reasoning reinforces the idea that our conclusions can be extended to other spacetimes without spherical symmetry, which means that our results also hold for other spacetime configurations, such as the state of the universe during the Planck epoch.
Finally, we emphasize that the scenario outlined above is general and model-independent. While alternative formulations of quantum gravity may predict slightly different values for some parameters, they cannot alter our conclusions: classical gravity gravitates, but quantum gravity antigravitates.
Acknowledgements.